# Entropy-Based Doppler Centroid Estimation and Speckle Noise Reduction for Spaceborne SAR Imaging

### v3

@@ II Model Description and Image Formation
where $f_{c}$ is the carrier frequency and the parameter $\beta$ is given as $b/T$, in which $b$ and $T$ stand for the bandwidth and the chirp time, respectively. In addition, $w_{r}$ is a rectangular window with length $T$ and $t$ is referred to as fast time. In addition, $\sigma$ is the complex reflectivity coefficient for the point reflector, $\eta$ is referred to as slow time and $R(\eta)$ is the distance from the radar to the target which is given as $R(\eta)=\sqrt{R^{2}_{0}+v^{2}(\eta-\eta_{c})^{2}}$. Moreover, $w_{a}$ is a rectangular window with its length equal to the synthetic aperture length divided by $v$, where $v$ is the speed of the satellite. The ultimate goal in SAR imaging is to estimate the complex reflectivity coefficient $\sigma$. The Range-Doppler algorithm is the first algorithm that was developed for space-borne SAR image reconstruction [10, 11, 12].
@@ III Doppler-Centroid Estimation
There are different techniques for the Doppler centroid frequency estimation and they are divided into two different categories, amplitude-based and phased-based methods [3]. The amplitude-based spectral fit algorithm and phase-based Average Cross Correlation Coefficient (ACCC) method are among two well-known methods for estimating $f^{\prime}_{dc}$ [3, 14]. The fractional part of the Doppler centroid frequency is needed to localize the energy of the targets int he azimuth direction. However, to compensate for the effect of RCM we need the unwrapped Doppler centroid frequency. The phased-based methods are sensitive to the noise and the amplitude-based techniques work well when there are isolated targets with strong reflectivity coefficients. Therefore, we propose the entropy-based approach as
@@
where $|\rm img[i,j]|$ represents the absolute value for the $(ij)^{\rm th}$ pixel of the reconstructed image. Furthermore, $m$ and $n$ are the number of pixels in the range and along track directions, respectively. Moreover, $\mathcal{I}$ is the entropy of the image.
@@ IV Speckle Noise
When the roughness of the terrain to be imaged is on the scale of the wavelength of the incident wave, speckle noise appears. Speckle noise is a multiplicative noise. In fact, the origin of the speckle noise is the dependency of the relative phase of the individual reflectors inside the resolution cell on the viewing angle [3].
The well-known method for speckle noise reduction is multi-look processing in which the synthetic aperture is divided into several independent sections and per each section an image from the scene is created [3]. The final image is the summation of the absolute value of the images created from the sub-apertures. In other words, non-coherent averaging of the reconstructed images yields the final image. The major problem with the multi-look processing is that the resolution reduces due to using only a small portion of the aperture when we create each independent images from the scene.
In this section, we utilize median filtering approach over the reconstructed image to reduce the effect of the speckle noise. We use the entire synthetic aperture to reconstruct the image, therefore, there is no loss in the resolution of the reconstructed image. We introduce a 2D $m\times n$ filter and slide it over the reconstructed image while solving the following optimization problem,
@@
Refer to captionAlong Track SamplesRange Samples(a)Refer to captionAmplitudeFrequency [Hz](b)
@@
Figure 5: The reconstructed images based on ${f^{\prime}}_{\rm dc}$ equal to a) $-1000\;\rm Hz$, b) $-500\;\rm Hz$, c) $0\;\rm Hz$, d) $500\;\rm Hz$, e) $516\;\rm Hz$, f) $1000\;\rm Hz$.
In order to present the entropy-based approach we select an isolated target with strong reflectivity coefficient in Fig. 4. In fact, we have shown one of the ships with a red circle in Fig. 4 and have used the data from this part of the scene to calculate the fractional part of the Doppler centroid. Fig. 5 illustrates the result of reconstructing the image based on different values for $f^{\prime}_{\rm dc}$. It is obvious from Fig. 5 that, when we are deviating from $520\;\rm Hz$, image de-focusing is occurred. To estimate the fractional part of the Doppler centroid using the entropy-based approach, we calculate the entropy of the reconstructed images using (8) and based on different values for $f^{\prime}_{\rm dc}$ ranging from $-1000\;\rm Hz$ to $1000\;\rm Hz$ with the step-size equal to $100\;\rm Hz$. The result has been presented in Fig. 6-(a). From Fig. 6-(a), it is clear that thee minimum entropy occurs in close proximity of $500\;\rm Hz$. Subsequently, to obtain a more accurate result we have searched through the range of frequencies from $400\;\rm Hz$ to $600\;\rm Hz$ and have shown the result in Fig. 6-(b). As it is clear from Fig. 6-(b) and specifically from the magnified part of the graph, the minimum value for the entropy is achieved when $f^{\prime}_{\rm dc}=516\;\rm Hz$. In fact, $f^{\prime}_{\rm dc}=516\;\rm Hz$ is the estimated value that we have obtained using the proposed entropy-based method.
@@
Figure 6: The entropy-based estimation of the fractional part of the Doppler centroid frequency for different values for $f^{\prime}_{\rm dc}$ ranging from a) $-1000\;\rm Hz$ to $1000\;\rm Hz$ with the step-size equal to $100\;\rm Hz$, b) $400\;\rm Hz$ to $600\;\rm Hz$ with the step-size equal to $10\;\rm Hz$ as well as the magnified part of the graph for $512\;\rm Hz$ to $520\;\rm Hz$ with the step-size equal to $1\;\rm Hz$.
Figure 7: The result of speckle noise reduction for the reconstructed image of English Bay shown in Fig. 4.
After the image reconstruction is completed, we will then focus on the speckle noise removal procedure based on (9). Fig. 7 shows the result of speckle noise reduction for the reconstructed image depicted in Fig. 4. To perform the speckle noise reduction, we have chosen $m=n=6$ for the filter given in (9).
Fig. 8 shows the result of applying the Range-Doppler algorithm to a larger part of the raw data which shows the entire Vancouver area. To remove the effect of the speckle noise we have applied the filter given in (9) with $m=n=10$.

### v4

@@ II Model Description and Image Formation
where $f_{c}$ is the carrier frequency and the parameter $\beta$ is given as $b/T$, in which $b$ and $T$ stand for the bandwidth and the chirp time, respectively. In addition, $w_{r}$ is a rectangular window with length $T$ and $t$ is referred to as fast time. In addition, $\sigma$ is the complex reflection coefficient for the point reflector, $\eta$ is referred to as slow time and $R(\eta)$ is the distance from the radar to the target which is given as $R(\eta)=\sqrt{R^{2}_{0}+v^{2}(\eta-\eta_{c})^{2}}$. Moreover, $w_{a}$ is a rectangular window with its length equal to the synthetic aperture length divided by $v$, where $v$ is the speed of the satellite. The ultimate goal in SAR imaging is to estimate the complex reflection coefficient $\sigma$. The Range-Doppler algorithm is the first algorithm that was developed for spaceborne SAR image reconstruction [10, 11, 12].
@@ III Doppler-Centroid Estimation
There are different techniques for the Doppler centroid frequency estimation and they are divided into two different categories, amplitude-based and phased-based methods [3]. The amplitude-based spectral fit algorithm and phase-based Average Cross Correlation Coefficient (ACCC) method are among two well-known methods for estimating $f^{\prime}_{dc}$ [3, 14]. The fractional part of the Doppler centroid frequency is needed to localize the energy of the targets in the azimuth direction. However, to compensate for the effect of RCM we need the unwrapped Doppler centroid frequency. The phased-based methods are sensitive to the noise and the amplitude-based techniques work well when there are isolated targets with strong reflection coefficients. Therefore, we propose the entropy-based approach as
@@
where $|\rm img[i,j]|$ represents the normalized absolute value for the $(ij)^{\rm th}$ pixel of the reconstructed image. Furthermore, $m$ and $n$ are the number of pixels in the range and along track directions, respectively. Moreover, $\mathcal{I}$ is the entropy of the image. When the image is focused, the entropy of the image is at its minimum level.
@@ IV Speckle Noise
When the roughness of the terrain to be imaged is on the scale of the wavelength of the incident wave, speckle noise appears. Speckle noise is a multiplicative noise. In fact, the origin of the speckle noise is the dependency of the relative phase of the individual reflectors inside the resolution cell on the viewing angle [3]. The well-known method for speckle noise reduction is multi-look processing in which the synthetic aperture is divided into several independent sections and per each section an image from the scene is created [3]. The final image is the summation of the absolute value of the images created from the sub-apertures. In other words, non-coherent averaging of the reconstructed images yields the final image. The major problem with the multi-look processing is that the resolution reduces due to using only a small portion of the aperture when we create each independent images from the scene. In this section, we utilize median filtering approach over the reconstructed image to reduce the effect of the speckle noise. We use the entire synthetic aperture to reconstruct the image, therefore, there is no loss in the resolution of the reconstructed image. We introduce a 2D $m\times n$ filter and slide it over the reconstructed image while solving the following optimization problem,
@@
[…]
@@
Figure 5: The reconstructed images of one of the ships, marked with red circle in Fig. 4, based on ${f^{\prime}}_{\rm dc}$ equal to, a) $-600\;\rm Hz$, b) $-300\;\rm Hz$, c) $0\;\rm Hz$, d) $300\;\rm Hz$, e) $516\;\rm Hz$, f) $600\;\rm Hz$.
In order to present the entropy-based approach we select an isolated target with strong reflection coefficient in Fig. 4. In fact, we have shown one of the ships with a red circle in Fig. 4 and have used the data from this part of the scene to calculate the fractional part of the Doppler centroid. Fig. 5 illustrates the result of reconstructing the image based on different values for $f^{\prime}_{\rm dc}$. It is obvious from Fig. 5 that, when we are deviating from $520\;\rm Hz$, image de-focusing is occurred. To estimate the fractional part of the Doppler centroid using the entropy-based approach, we calculate the entropy of the reconstructed images using (8) and based on different values for $f^{\prime}_{\rm dc}$ ranging from $-600\;\rm Hz$ to $600\;\rm Hz$ with the step-size equal to $100\;\rm Hz$. The result has been presented in Fig. 6-(a). From Fig. 6-(a), it is clear that thee minimum entropy occurs in close proximity of $500\;\rm Hz$. Subsequently, to obtain a more accurate result we have searched through the range of frequencies from $400\;\rm Hz$ to $600\;\rm Hz$ and have shown the result in Fig. 6-(b). As it is clear from Fig. 6-(b) and specifically from the magnified part of the graph, the minimum value for the entropy is achieved when $f^{\prime}_{\rm dc}=516\;\rm Hz$. In fact, $f^{\prime}_{\rm dc}=516\;\rm Hz$ is the estimated value that we have obtained using the proposed entropy-based method.
@@
Figure 6: The entropy-based estimation of the fractional part of the Doppler centroid frequency for different values for $f^{\prime}_{\rm dc}$ ranging from a) $-600\;\rm Hz$ to $600\;\rm Hz$ with the step-size equal to $100\;\rm Hz$, b) $400\;\rm Hz$ to $600\;\rm Hz$ with the step-size equal to $10\;\rm Hz$ as well as the magnified part of the graph for $512\;\rm Hz$ to $520\;\rm Hz$ with the step-size equal to $1\;\rm Hz$.
Figure 7: The result of the speckle noise reduction process, based on (9), for the reconstructed image of English Bay shown in Fig. 4.
After the image reconstruction is completed, we will then focus on the speckle noise removal procedure based on (9). Fig. 7 shows the result of speckle noise reduction for the reconstructed image depicted in Fig. 4. To perform the speckle noise reduction, we have chosen $m=n=6$ for the filter given in (9). Fig. 8 shows the result of applying the Range-Doppler algorithm to a larger part of the raw data which shows the entire Vancouver area. To remove the effect of the speckle noise we have applied the filter given in (9) with $m=n=10$.